OpenStudy (anonymous):

Solve by Substitution 2x-y=6 x+y=-3

OpenStudy (anonymous):

You can choose from 3 methods (substitution, graphing, and elimination). But they already have elimination set up for you. Notice the positive y and the negative y. line up the two equations and the y variables cancel out and figure out X. 2x - y = 6 x +y =-3 3x = 3 x=1 then substitute: x+y =-3 1+y=-3 y=-4

OpenStudy (anonymous):

Okay, thank you!